These reaction times (in tenths of a second) were recorded for group of subjects after each had been given in a drug pain.


Drug A Drug B Drug C

4 9 8

7 11 6

6 12 7

3 8 6

4 10 5

3 11 7


1. Formulate the H0, and H1:

2. C.V. at alpha .01 level of significance

3. Compute the test value: ANOVA - F ratio

4. Decision

5. Interpretation


1
Expert's answer
2020-07-14T18:04:31-0400

The total sample size is N=18.


N=18. Therefore, the total degrees of freedom are:


dftotal=18-1=17


Also, the between-groups degrees of freedom are dfbetween=3-1=2, and the within-groups degrees of freedom are:


dfwithin=dftotal-dfbetween=17-2=15

"\displaystyle\sum_{i,j}x_{ij}=27+61+39=127"

"\displaystyle\sum_{i,j}x_{ij}^2 =135+631+259=1025"

"SS_{total}=\displaystyle\sum_{i,j}x_{ij}^2-{1\over N}(\displaystyle\sum_{i,j}x_{ij})^2=1025-{127^2\over 18}=128.944"

"SS_{within}=\displaystyle\sum SS_{within\ groups}
=13.5+10.833+5.5=29.833"

"MS_{between}={SS_{between}\over df_{between}}={128.944-29.833\over 2}=49.556; 
MS_{within}={SS_{within}\over df_{within}}={29.833\over 15}=1.989"

"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{49.556}{1.989}=24.916"

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2=\mu_3"

H1: Not all means are equal

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.

Based on the information provided, the significance level is "\alpha" =0.01,

 and the degrees of freedom are df1=2, df2=2, therefore, the rejection region for this F-test is R="\{F: F>F_C=6.359\}"


Since it is observed that F=24.916>6.359=FC,

​, it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that not all 3 population means are equal, at the   α =0.01

  significance level.

Using the P-value approach: The p-value is  p=0, using table and since p=0<0.01, it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that not all 3 population means are equal, at the  α =0.01significance level.